An investment of R20 000 accumulated to R45 200. If the applicable simple interest rate is 12% per year, then the time under consideration is

[1] 10,50 years.

[2] 15,12 years.

[3] 7,19 years.

[4] 4,65 years.

[5] 3,25 years


1
Expert's answer
2021-06-15T15:01:26-0400

Given,

P= R 20,000

A = R45200

r=12%

 

let time = t

"A=P×(1+R×T)\\45200=20000×(1+0.12×T)\\T=\frac{(\frac{45200}{20000}−1)}{0.12}\\=10.50 \space YEARS"